### v2

# FireANTs: Adaptive Riemannian Optimization for Multi-Scale Diffeomorphic Matching

Abstract
The paper proposes FireANTs, the first multi-scale Adaptive Riemannian Optimization algorithm for dense diffeomorphic image matching. One of the most critical and understudied aspects of diffeomorphic image matching algorithms are its highly ill-conditioned nature. We quantitatively capture the extent of ill-conditioning in a typical MRI matching task, motivating the need for an adaptive optimization algorithm for diffeomorphic matching. To this end, FireANTs generalizes the concept of momentum and adaptive estimates of the Hessian to mitigate this ill-conditioning in the non-Euclidean space of diffeomorphisms. Unlike common non-Euclidean manifolds, we also formalize considerations for multi-scale optimization of diffeomorphisms. Our rigorous mathematical results and operational contributions lead to a state-of-the-art dense matching algorithm that can be applied to generic image data with remarkable accuracy and robustness. We demonstrate consistent improvements in image matching performance across a spectrum of community-standard medical and biological correspondence matching challenges spanning a wide variety of image modalities, anatomies, resolutions, acquisition protocols, and preprocessing pipelines. This improvement is supplemented by from 300$\times$ up to 3200$\times$ speedup over existing state-of-the-art algorithms. For the first time, we perform diffeomorphic matching of sub-micron mouse cortex volumes at native resolution. Our fast implementation also enables hyperparameter studies that were intractable with existing correspondence matching algorithms.
Keywords: correspondence matching, deformable image matching, diffeomorphisms, optimization, non-Euclidean manifold, microscopy, neuroimaging
1 Introduction
The ability to identify and map corresponding elements across diverse datasets or perceptual inputs – known as correspondence matching – is fundamental to interpreting and interacting with the world. Correspondence matching between images is one of the longstanding fundamental problem in computer vision. Influential computer vision researcher Takeo Kanade famously once said that the three fundamental problems of computer vision are: “Correspondence, correspondence, correspondence” 1. Indeed, correspondence matching is fundamental and ubiquitous across various disciplines, manifesting in many forms including but not limited to stereo matching  2, structure from motion  3, 4, template matching  5, motion tracking  6, 7, shape correspondence  8, semantic correspondence  9, point cloud matching  10, optical flow  11, and deformable image matching  12. Solving these problems addresses the desiderata for a wide range of applications in computer vision, robotics, medical imaging, remote sensing, photogrammetry, geological and ecological sciences, cognitive sciences, human-computer interaction, self-driving among many other fields.
Correspondence matching is broadly divided into two categories: sparse and dense matching. Most sparse matching problems like stereo matching, structure from motion, and template matching involve finding a sparse set of salient features across images followed by matching them. In such cases, the transformation between images, surfaces, or point clouds is typically also parameterized with a small number of parameters, e.g., an affine transform, homography or a fundamental matrix. These methods are often robust to noise, occlusions, and salient features can be detected and matched efficiently via analytical closed forms. In contrast, dense matching is much harder because the entire image is considered for matching and cannot be reduced to a sparse set of salient features, and the transformation between images is typically parameterized with a large number of parameters, e.g., a dense deformation field. Moreover, dense matching is sensitive to local noise, and cannot be solved efficiently via analytical closed forms – necessitating iterative optimization methods 13, 14, 15, 16, 17, 18. Due to the dense and high-dimensional nature, these methods are often plagued with ill-posedness, curse of dimensionality, difficulty in optimization and scalability to high-resolution data.
In this work, we focus on dense deformable correspondence matching, which is the non-linear and local (hence deformable) alignment of two or more images into a common coordinate system. Dense deformable correspondence matching is a fundamental problem in computer vision 19, medical imaging 20, 21, 22, microscopy 23, 24, and remote sensing. Here, we focus on applications in biomedical and biological imaging. In the biomedical and biological sciences, deformable correspondence matching is also referred to as deformable registration. Within dense deformations, diffeomorphisms are of special interest as a family of deformations that are invertible transformations such that both the transform and its inverse are differentiable. This allows us to accurately model the deformation between images and also ensures that the deformation is physically plausible.
We address and tackle two fundamental problems in dense correspondence matching: ill-conditioning and scalability. The ill-conditioning arises due to the high-dimensional and heterogeneous nature of the dense matching optimization objective, that can be mitigated by adaptive optimization methods. Although standard adaptive optimization methods 25, 26 are shown to work in fixed Euclidean spaces, it is not obvious how to extend this formulation to the non-Euclidean space of diffeomorphisms. Fortunately, diffeomorphisms admit many interesting mathematical properties like being embedded in a Riemannanian manifold, having a Lie Group structure, and local geodesic formulations that can be exploited for adaptive optimization. We present a novel and mathematically rigorous framework for adaptive optimization of diffeomorphic matching Section 4.4. This is done by exploiting the group structure of diffeomorphisms to define a custom gradient descent algorithm, followed by adaptive optimization on this space. Second, we observe that most existing state-of-the-art methods are prohibitively slow for high-resolution data, which limits their applicability to rigorous hyperparameter studies, large-scale data, or high-resolution alignment at mesoscopic or microscopic resolutions. Our novel operational contributions lead to an algorithm that is 3200$\times$ faster than existing state-of-the-art toolkits (Fig. 4) and can perform diffeomorphic matching on modalities like microscopy images at native resolution (Fig. 3). This unprecedented speedup and scalability makes large-scale hyperparameter studies computationally feasible (Figs. 5 and 4). We package our contributions into a software toolkit called FireANTs, which is an open-source state-of-the-art toolkit for dense deformable correspondence matching. Our framework can allow a practitioner to perform interactive dense matching that is useful for aligning complex, multimodal, multi-channel or high-resolution images 27, 28 or to provide guidance to compensate for missing data (e.g., in microscopy imaging), wherein a typical dense image matching algorithm would fail to run interactively.
2 Results
FireANTs represents the next generation of algorithms superseding the widely established and successful adoption of the ANTs ecosystem spanning the gamut of biomedical and life sciences research. Accordingly, we rigorously evaluate FireANTs against established benchmarks to substantiate its efficacy and its viability as a robust registration algorithm across anatomies and modalities. Furthermore, we extend our analysis to encompass datasets with considerably higher image heterogeneity, namely deformable registration of a high-resolution expansion microscopy benchmark at native resolution. Finally, we demonstrate its unprecedented speedup and usability by showing speedups of up to three orders of magnitude and robustness to hyperparameters, subsequently opening avenues in the ability to perform large-scale hyperparameter studies and registration of exceedingly complex and larger datasets.
2.1 Experiment Setup
We briefly describe the significance, existing state-of-the-art and challenges associated with the chosen benchmarks to demonstrate the efficacy of FireANTs.
Klein et al. brain mapping challenge  29
Within the biomedical sciences, registration algorithms like ANTs have been predominantly used by the neuroimaging community by far. Klein et al. 29 in their landmark paper reported an extensive evaluation of fourteen state-of-the-art registration algorithms on four neuroimaging datasets. The four neuroimaging datasets (IBSR18, CUMC12, MGH10, LPBA40) comprise different whole-brain labelling protocols, eight different evaluation measures and three independent analysis methods. More details about the datasets are outlined in  Section A.1. ANTs was one of the top performing methods for this challenge, performing well robustly across all four datasets. Evaluating on this challenge is therefore imperative to establish FireANTs as an effective, versatile and robust algorithm for neuroimaging applications.
A natural way to evaluate whether two images are in a common coordinate frame is to evaluate the accuracy of overlap of gross morphological anatomical structures. The method considers measures of volume and surface overlap, volume similarity, and distance measures to evaluate the alignment of anatomical regions. Given a source label map ${S}_{r}$ and target label map ${T}_{r}$ and a cardinality operator |.||.|| . |, we consider the following overlap measures. The first measure ‘target overlap’, defined as the overlap between the source and target divided by the target. Target overlap is a measure of sensitivity, and the original evaluation 29 considers the aggregate total overlap as follows
$\small TO_{r}=\frac{|{S}_{r}\cap{T}_{r}|}{|{T}_{r}|},TO_{Klein}=\frac{\sum_{r}|{S}_{r}\cap{T}_{r}|}{\sum_{r}|{T}_{r}|}$ (1)
However, we notice that this measure of overlap is biased towards larger anatomical structures, since both the numerator $\sum_{r}|{S}_{r}\cap{T}_{r}|$ and denominator $\sum_{r}|{T}_{r}|$ sums are dominated by regions with larger number of pixels. To normalize for this bias, we also consider a target overlap that is simply the average of region-wise target overlap.
$\small TO=\frac{1}{N_{r}}\sum_{r}TO_{r}$ (2)
We also consider a second measure, called mean overlap (MO), more popularly known as the Dice coefficient or Dice score. It is defined as the intersection over mean of the two volumes. Similar to target overlap, we consider two aggregates of the mean overlap over regions:
$MO_{r}=2\frac{|{S}_{r}\cap{T}_{r}|}{|{S}_{r}|+|{T}_{r}|},MO_{Klein}=2\frac{\sum_{r}|{S}_{r}\cap{T}_{r}|}{\sum_{r}(|{S}_{r}|+|{T}_{r}|)},MO=\frac{1}{N_{r}}\sum_{r}MO_{r}$ (3)
Klein et al.29 also propose a ‘Union Overlap’ metric which is a monotonic function of the Dice score. Therefore, we do not use this in our evaluation. To complement the above agreement measures, we also compute false negatives (FN), false positives (FP), and volume similarity (VS) coefficient for anatomical region $r$:
$FN_{r}=\frac{|{T}_{r}\backslash{S}_{r}|}{|{T}_{r}|},\quad FP_{r}=\frac{|{S}_{r}\backslash{T}_{r}|}{|{S}_{r}|},\quad VS_{r}=2\frac{|{S}_{r}|-|{T}_{r}|}{|{S}_{r}|+|{T}_{r}|}$ (4)
Similar to the overlap metrics, we compute the aggregates as in the original evaluation denoted by $FN_{Klein},FP_{Klein},VS_{Klein}$ and average over regions denoted simply by FN, FP, VS. This leads to a total of 10 aggregate metrics that we use to compare our method with 4 baselines - ANTs, Demons 18, VoxelMorph 30 and SynthMorph 31, representing established classical and deep learning registration algorithms.
EMPIRE10 challenge  22
Alignment of thoracic CT images, especially the lung and its internal structures is a challenging task, owing to the highly deformable nature of the lungs. Pulmonary registration has extensive clinical utility, such as aligning temporally distinct breath-hold scans to facilitate visual comparison and reduce errors. It also aids in modeling lung expansion by registering inspiration and expiration scans, and tracking disease progression by aligning breath-hold scans over time to enhance monitoring of airflow and pulmonary function. Murphy et al. propose the Evaluation of Methods for Pulmonary Image REgistration 2010 (EMPIRE10) challenge to provide a platform for a comprehensive evaluation and fair comparison of registration algorithms for the task of CT lung registration. The dataset consists of 30 scan pairs including inspiration-expiration scans, breathhold scans over time, scans from 4D data, ovine data, contrast-noncontrast, and artificially warped scan pairs. The challenge provides a broad range of data complexity, voxel sizes and image acquisition differences, and the challenge still continues as of today. The challenge only provides scan pairs and binary lung masks. All the other data (fissures and landmarks) are withheld for evaluation. All the scan pairs have varying spatial and physical resolutions, are acquired over a varying set of imaging configurations. This calls for a registration algorithm that is agnostic to any assumptions about anisotropy of image resolution, both physical and voxel. We use the evaluation provided by the challenge, and compare the fissure alignment, landmark alignment, and singularity of registration. More details about the evaluation can be found in  22. We compare our method with ANTs which performs direct gradient descent updates and DARTEL which optimizes a stationary velocity field using the metrics reported in the evaluation server.
RnR ExM mouse dataset  32
[…]
Figure 1: FireANTs demonstrates state-of-the-art performance on Klein et al.challenge 29: Following the evaluation setup of Klein et al. paper, we validate registration performance using the average volume overlap of all anatomical label maps between the fixed and warped label maps. We consider ANTs (the winner of the challenge), and Diffeomorphic Demons as state-of-the-art optimization algorithms, and Voxelmorph and Synthmorph as state-of-the-art unsupervised deep learning baselines. Evaluation is shown for five metrics with $\uparrow$ denoting a higher score is better, and $\downarrow$ signifying a lower score is better. For deep learning baselines, appropriate preprocessing (intensity normalization, alignment, and resampling to 1mm isotropic) is performed to ensure a fair comparison, whereas no such preprocessing is required for optimization methods, including FireANTs. FireANTs shows significant gains in performance that are consistent across all four datasets, with the median overlap scores outperforming the third quartile of all other methods for IBSR18 and CUMC12 datasets. Comparison of overlap metrics by specific anatomical regions are in Fig. 10. For the overlap aggregation mentioned in 29, results are shown in Fig. 11.
2.2 Results on state-of-the-art biomedical benchmarks
Klein et al. human brain MRI registration
Klein et al. 29 present a landmark challenge with comprehensive, unbiased, and thorough evaluation of different registration algorithms on four MRI brain datasets, with Advanced Normalization Tools (ANTs) 13 being the top performing method overall. We compare our method with two state-of-the-art optimization algorithms: ANTs - which won the original challenge, and Symmetric Log Demons 18, and two widely used deep learning algorithms: VoxelMorph 30 and SynthMorph 31 using their provided pretrained models. For all the four datasets, we first fit an affine transformation from the moving image to the fixed image, followed by a diffeomorphic transform. Results for the brain datasets are shown in  Fig. 1 and  Fig. 10.
Our algorithm outperforms all baselines on all four datasets, with an improvement in all metrics evaluating the volume overlap of the fixed and warped label maps. The improvements are consistent in all datasets, with varying number and sizes of anatomical label maps. In the IBSR18 and CUMC12 datasets, the median target overlap of our method is better than the third-quantile of ANTs.  Fig. 10 also highlights the improvement in label overlap per labeled brain region across all datasets. A caveat with deep learning methods is that their performance is highly dependent on the domain gap between the training and test datasets. VoxelMorph is trained on the OASIS dataset, which has different image statistics compared to the four datasets, and consequently we see a performance drop. Moreover, VoxelMorph is sensitive to the anisotropy of the volumes, consequently all volumes are resampled to 1mm isotropic, and renormalized for VoxelMorph. A noticable performance drop is observed when the anisotropic volumes are fed into the network, which is undesirable as the trained model is essentially ‘locked’ to a single physical resolution - which limits the generalizability of the model to various modalities with different physical resolutions. For Demons, ANTs, and FireANTs (Ours), we do not perform any additional normalization or resampling. SynthMorph is more robust to the domain gap than VoxelMorph due to its training strategy with synthetic images, but still underperforms optimization baselines when their recommended hyperparameters are chosen.
(a) FireANTs achieves substantially lower inter-quartile range of fissure errors, defined as the percentage of marked pixels that are registered to points on the opposite side of the fissure boundary.
% Error DARTEL ANTs Ours
Left Lung 3.9983 0.0069 0.0000
Lower Lung 2.7514 0.0177 0.0000
Right Lung 2.4930 0.0107 0.0000
Upper Lung 5.2037 0.0000 0.0000
Score (Overall) 3.0681 0.0088 0.0000
(b) Singularity errors are defined as fraction of voxels that define a non-invertible deformation. Singularity quantifies the percentage of implausible deformations. FireANTs achieves zero percent singularity errors.
(c) Landmark distance is the Euclidean distance between well-dispersed landmark points between the fixed and warped images. FireANTs has a lower median and narrower interquartile range than baselines on five out of six subregions.
Method Left Lung Right Lung Score
(Error Overall)
FireANTs (Ours) 0.0185 0.0254 0.0227
MRF Correspondence Fields 0.0824 0.0211 0.0485
ANTs 0.0249 0.1016 0.0747
Dense Displacement Sampling 0.0578 0.0919 0.0826
ANTs + BSpline 0.0821 0.0848 0.0861
DISCO 0.1256 0.0499 0.0882
VIRNet 0.0834 0.0934 0.0890
Feature-constrained nonlinear registration 0.1210 0.0758 0.1032
Explicit Boundary Alignment 0.1063 0.1246 0.1209
MetaReg 0.1049 0.2224 0.1791
(d) Fissure alignment error on top 10 algorithms in the challenge sorted by fissure alignment error, averaged on all scan pairs. FireANTs outperforms a wide array of baselines, including direct optimization (ANTs, ANTs+BSpline), neural networks (VIRNet), and explicit correlation volumes (MRF, Disco).
Figure 2: FireANTs demonstrates state-of-the-art performance on EMPIRE10 lung registration challenge: (a): Lung fissure plates are an important anatomical landmark demarcating lobes within the lung. Fissure alignment errors (in %) denote the percentage of locations near the lung fissure plates that are on the wrong side of the fissure post-registration. Over all 30 scan pairs, our method performs 5$\times$ better than ANTs. (b): Singularity errors defined as percentage of voxels that have a non-diffeomorphic deformation, a proxy for physically implausible deformations. In the DARTEL baseline, singularities can be introduced for larger deformations due to numerical approximations of the integration. Even for ANTs, the solutions (deformations) returned are not entirely diffeomorphic. Our method shows much better fissure and landmark alignment (Fig. 2(a,c),  Fig. 13Fig. 14) with guaranteed diffeomorphic transforms. (c): Landmark distance in mm for selected landmarks. Across different lung subregions, our method shows results at least at par with ANTs, with slightly better average and median results across all regions. (d): Shows the top 10 algorithms for average fissure alignment error in % in the EMPIRE10 challenge. Error metrics are taken from the evaluation server. Other methods perform well on one lung (MRF for right, ANTs for left) but comparatively poorly on the other lung, compared to our method showing both accurate and robustness to both the left and right lung.    X = First,    X = Second,    X = Third best result.
EMPIRE10 lung CT challenge
The EMPIRE10 challenge  22 is an established benchmark challenge and it provides a platform for in-depth evaluation and fair comparison of available registration algorithms for this application. We discuss more details about the challenge in  Section 2.1. The lung images represent volumes that are about 10$\times$ larger than the brain dataset, thereby presenting a significant challenge for deformable registration algorithms. ANTs is, again one of the top performing methods in this challenge. Unlike the brain datasets, ground truth labels for fissure and landmarks are not provided for validation. Therefore, we rely on the evaluation metrics computed in the evaluation server. We compare our method with two powerful baselines (i) ANTs, which optimizes the diffeomorphism directly, and (ii) the DARTEL 17 formulation optimizing a stationary velocity field (SVF), where the diffeomorphism is obtained using an exponential map of the SVF. We first affinely align the binary lung masks of the moving and fixed images using Dice loss 33. This is followed by a diffeomorphic registration using the intensity images.
We consider three evaluation criteria of the challenge - (1) fissure alignment errors (in %) denoting the fraction of fissure voxels that are misaligned after registration (Figs. 2 and 2), (2) landmark distance (in mm) (Fig. 2), and (3) singularity errors - which is defined as the fraction of the image volume that is warped non-diffeomorphically (Fig. 2). Results in  Fig. 2 also demonstrates the effect of representation choice for modeling diffeomorphisms. For the same scan pairs and cost functions, the exponential map-based DARTEL performs substantially poorly in terms of fissure alignment, landmark distance and singularities than that of ANTs by three orders of magnitude. In contrast, our method has about a 5$\times$ lower error than ANTs on the fissure alignment task, and performs better on 5 out of 6 subregions on the landmark distance alignment task. Moreover, although all methods return deformations that are theoretically diffeomorphic, the SVF representation introduces significant singularity errors (voxels where the deformation is not diffeomorphic) due to non-adaptive value of $N$ in the scaling-and-squaring step. The ANTs baseline also introduces some singularities in its proposed diffeomorphic transform. Our method, on the other hand computes numerically perfect diffeomorphic transforms. Finally, we compare the fissure alignment error of all submissions in the EMPIRE10 challenge, and show the top 10 algorithms in  Fig. 2. Results demonstrate that FireANTs attains the lowest landmark alignment errors compared to an array of contemporary state-of-the-art algorithms. FireANTs also has the lowest runtime of all 10 methods, signifying a paradigm shift in the balance between computational efficiency and accuracy instigated by our approach.
2.3 Evaluation on high-resolution mouse cortex registration
[…]
In contrast, FireANTs is able to register the volume at native resolution. We perform an affine registration followed by a diffeomorphic registration step. The entire method takes about 2-3 minutes on a single A6000 GPU. As shown in  Fig. 3, our method secures the first place on the leaderboard, with a considerable improvement in the Dice score and a 4.42$\times$ reduction in the standard deviation of the Dice scores compared to the next best method.  Fig. 3 also shows qualitative comparison of our method compared to Bigstream 38, the winner of the RnR-ExM challenge. Bigstream only performs an affine registration, leading to inaccurate registration in one of three test volumes, leading to a lower average Dice score and higher variance. Moreover, the affine registration leads to boundary in-plane slices being knocked out of the volume, leading to poor registration (Fig. 3). FireANTs preserves the boundary in-plane slices during its affine step, and subsequently performs an accurate diffeomorphic registration. This shows the versatility and applicability of FireANTs for high-resolution microscopy registration.
(a) Snapshot of the RnR-ExM leaderboard
(b) Qualitative comparison of registration of Bigstream and FireANTs
Figure 3: FireANTs secures first rank in the RnR-ExM mouse dataset: (a): As of January 17, 2025, our method ranks first in the mouse brain registration task, which is the only task in the challenge requiring deformable registration. Our top two successful submissions secure the first and second position, with a 0.361 improvement in Dice score compared to the 3rd ranked submission, which is 0.261 better than the 5th ranked submission (bigstream). Note that among the top 10 submissions, our method has the lowest standard deviation (4.42$\times$ lower than the second best submission) showing the robustness of our model across different microscopy volumes. (b) shows a qualitative comparison of FireANTs with Bigstream  38, the other top leading method in the challenge. The moving image volumes have substantially more noise than the fixed image volumes, making intensity-based registration difficult. The non-rigid deformation dynamics of the hydrogel are clearly visible, as the moving volume has a thicker boundary than the fixed volume. Bigstream does not capture these dynamics very well – this is illustrated by comparing the thickness of the cortex at various points (zoomed orange crops in bottom row), where Bigstream does not deform the cortex enough to match the fixed image. FireANTs deforms and accurately depicts these morphological changes, which can be crucial for downstream morphometric studies. Moreover, the affine registration in Bigstream knocks the boundary slices out of the volume (red highlight in top row), leading to drop in registration performance. On contrary, our method’s affine and deformable stages are more stable, leading to better registration and avoiding spurious out-of-bound artifacts at the boundary slices.
2.4 Ease of experimentation due to efficient implementation
One of the major contributions of our work is to enable fast and scalable image registration while improving accuracy. In applications like atlas/template building, registration is used in an iterative manner (in the ‘inner loop’) of the optimization. Another application that requires fast runtimes is hyperparameter tuning, since different datasets and modalities admit notably different hyperparameters for optimal registration. This calls for an rapidly escalating need for fast and scalable registration algorithms. To demonstrate the computational and runtime efficiency of our method, we demonstrate the runtime of our library on the brain and lung benchmarks. All the experiments for our method are run on a single A6000 GPU, with a batch size of 1, to avoid amortizing runtime over a bigger batch size. In practice, higher speedups can be obtained by registering batches of images that fit in a single GPU. For the brain datasets, we run ANTs with the recommended configuration with AMD EPYC 7713 Processor (single thread) and 512GB RAM. For the EMPIRE10 lung dataset, we use the runtimes described in the writeup provided as part of the challenge. A runtime analysis of our method on the brain and EMPIRE10 datasets are shown in Fig. 4.
(a) Histogram and summary of runtime on four brain MRI datasets.
ANTs Ours Speedup
Avg 10380.57 6.11 1701.01
Min 8645.76 5.87 1424.66
Max 11438.96 6.99 1930.78
ANTs Ours Speedup
Avg 15497.14 6.30 2475.39
Min 13298.00 5.79 2002.77
Max 18983.44 6.80 3270.91
ANTs Ours Speedup
Avg 10558.51 4.25 2486.57
Min 5796.64 3.66 1187.98
Max 15524.08 4.88 3773.99
ANTs Ours Speedup
Avg 7273.74 7.02 1036.03
Min 5692.80 6.36 821.20
Max 12068.08 7.66 1635.29
(b) Runtime analysis and summary on EMPIRE10 dataset
ANTs DARTEL Ours Speedup (ANTs) Speedup (DARTEL)
Avg 6hr 14m 7hr 16m 0m 39s 562.67 663.77
Min 0h 55m 1h 8m 0m 9s 320.74 315.23
Max 12h 41m 10h 11m 1m 5s 1231.27 796.51
Figure 4: FireANTs runs with unprecedented speedups. We compare the runtime of our implementation with the ANTs library. (a) shows histogram of speedup (runtime of ANTs divided by runtime of our method) and statistics of runtimes (in seconds) for the four brain MRI datasets. For all datasets, our implementation runs a minimum of two orders of magnitudes faster, making it suitable for hyperparameter search algorithms, and larger datasets. Table (b) shows the runtime of ANTs, DARTEL and our implementation on the EMPIRE10 challenge data. The first three colums show the actual runtime of the methods, followed by the speedup obtained by our method when compared to ANTs and DARTEL. Note that our method runs a minimum of 320 times faster than ANTs, saving a substantial amount of time, at no loss in registration quality.
(a) Grid search on LPBA40 dataset
(b) Grid search on the EMPIRE10 dataset, with Dice score
Figure 5: FireANTs facilitates feasibilility of extensive hyperparameter search in registration The speed of FireANTs makes hyperparameter studies like these feasible, which ANTs would take years to complete. (a): We perform a hyperparameter grid search on three hyperparameters of interest - smoothing kernel for the warp field ($\sigma_{\text{warp}}$) in pixels, smoothing kernel for the gradient of warp field ($\sigma_{\text{grad}}$) in pixels and learning rate $\eta$. The metric to optimize in this case is the target overlap. For the LPBA40 dataset, we perform a hyperparameter sweep over 640 configurations in 40 hours with 8 A6000 GPUs. A corresponding hyperparameter sweep with 8 concurrent jobs with each job consuming 8 CPUs would take $\sim$3.6 years to complete. The white contour representing the level set for target overlap = 0.75, and the black contour representing the level set for target overlap of 0.74 show the robustness of our method to hyperparameters - performance is not brittle or sensitive to choice of hyperparameters. (b): Hyperparameter grid search on the EMPIRE10 dataset over $\sigma_{\text{warp}}$ and $\sigma_{\text{grad}}$ parameters (456 configurations), with a fixed learning rate of $\eta=0.25$. The metric to optimize is the Dice score of the provided binary lung mask. This sweep takes about 12.37 hours on 8 GPUs, whereas a corresponding sweep would take 296 days for ANTs and 345 days for DARTEL (more in Fig. 4). The white contour corresponds to the level set for Dice score = 0.96, showing both a huge spectrum of parameters that achieve high Dice scores, and low sensitivity of the method to choice of hyperparameters.
FireANTs shows up to three orders faster runtime than ANTs
For the EMPIRE10 dataset, we compare our method with both ANTs and DARTEL implementations. We capture the runtime for each scan pair and aggregate the average, minimum and maximum runtimes and speedups in Fig. 4. FireANTs reduces the runtime from 1 to 12 hours for a single scan pair to under a minute. Since the exponential map requires a few integration steps for each iteration and pullback of the gradient from the warp to the velocity field, this variant is significantly slower than ANTs. In contrast, our method achieves a minimum of more than 300$\times$ speedup over ANTs, while maintaining high accuracy. In the brain datasets, our method achieves a consistent speedup of three orders of magnitude. This occurs due to a better choice of hyperparameters compared to the baseline, faster convergence due to the adaptive optimization, and better memory and compute utilization by efficient cuDNN implementations.
On the brain datasets, we compare our method with ANTs, and compute the runtime and speedups for each scan pair. We then plot a histogram of the speedup and aggregate the runtime and speedup in Fig. 4. On all datasets, the average speedup is over 1000$\times$, signifying a new paradigm for image registration workflows. These improvements in runtime occur while also providing at par, or superior quantitative and qualitative results (Fig. 12, Extended Data 1314).
FireANTs allows fast hyperparameter tuning
In optimization toolkits such as ANTs, correct choice of hyperparameters are key to high quality registration. Some of these hyperparameters are the window size for the similarity metric Cross-Correlation or bin size for Mutual Information. In our experience, the Gaussian smoothing kernel $\sigma_{\text{grad}},\sigma_{\text{warp}}$ for the gradient and the warp field are two of the most important parameters for accurate diffeomorphic registration. The optimal values of these hyperparameters vary by image modality, intensity profile, noise and resolution. Typically, these values are provided by some combination of expertise of domain experts and trial-and-error. However, non experts may not be able to adopt these parameters in different domains or novel acquisition settings. Recently, techniques such as hyperparameter tuning have become popular, especially in deep learning.
In the case of registration, hyperparameter search can be performed by considering some form of label/landmark overlap measure between images in a validation set. We demonstrate the stability and runtime efficiency of our method using two experiments : (1) Owing to the fast runtimes of our implementation, we show that hyperparameter tuning is now feasible for different datasets. The optimal set of hyperparameters is dependent on the dataset and image statistics, as shown in the LPBA40 and EMPIRE10 datasets; (2) within a particular dataset, the sensitivity of our method around the optimal hyperparameters is very low, demonstrating the robustness and reliability of our method. We choose the LPBA40 dataset among the 4 brain datasets due to its larger size (40$\times$39 = 1560 pairs). We choose three parameters to search over : the learning rate ($\eta$), and the gaussian smoothing parameters $\sigma_{\text{warp}},\sigma_{\text{grad}}$. We use the Ray library (https://docs.ray.io/) to perform a hyperparameter tuning using grid search. For the LPBA40 dataset, a grid search over three parameters (shown in  Fig. 5) takes about 40.4 hours with 8 parallel jobs. On the contrary, ANTs would require around 3.6 years to complete the same grid search, with 8 threads allocated to each job and 8 parallel jobs. This makes hyperparameter search for an unknown modality computationally tractable. Fig. 5 shows a dense red region suggesting the final target overlap is not sensitive to the choice of hyperparameters. Specifically, the maximum target overlap is 0.7565 and 58.4% of these configurations have an average target overlap of $\geq$ 0.74. This is demonstrated in Fig. 5 by the white contour line denoting the level set for target overlap = 0.75, and the black contour line denoting the level set for target overlap of 0.74. The target overlap is quite insensitive to the learning rate ($\geq 0.4$) showing that our algorithm achieves fast convergence with a smaller learning rate. On the EMPIRE10 dataset, we fix the learning rate and perform a similar hyperparameter search over two parameters, the Gaussian smoothing parameters $\sigma_{\text{warp}},\sigma_{\text{grad}}$, shown in Fig. 5. We use the average Dice score between the fixed and moving lung mask to choose the optimal hyperparameters. FireANTs can perform a full grid search over 456 configurations on the EMPIRE10 dataset in 12.37 hours with 8 A6000 GPUs, while it takes SyN 10.031 days to run over a single configuration. Normalizing for 8 concurrent jobs and 456 configurations, it would take ANTs about 296 days, and DARTEL about 345 days. This shows that our method and accompanying implementation can now make hyperparameter search for 3D image registration studies feasible.
3 Discussion
We present FireANTs, a powerful and general-purpose multi-scale registration algorithm. Our method performs registration by generalizing the concept of first-order adaptive optimization schemes for optimizing parameters in a fixed Euclidean space, to multi-scale diffeomorphisms. This generalization is highly non-trivial because diffeomorphisms are typically implemented as an image grid proportional to the size of the fixed image, and are optimized in a multi-scale manner to capture large deformations 40, 13, 17 leading to changing grid size throughout optimization. Our method also avoids computationally expensive parallel transport and riemannian metric tensor computation steps for diffeomorphisms by solving an Eulerian descent that exploits the group structure to define descent directions from the identity transform. FireANTs achieves consistent improvements in performance over state-of-the-art optimization-based registration algorithms like ANTs, DARTEL, SynthMorph and Bigstream. This improvement is shown across six datasets with a spectrum of anatomies (in-vivo human brain, human, ovine and synthetic lungs, mouse cortex), contrast, image volume sizes (ranging from 196 up to 2048 voxels per dimension), and modalities (MRI, CT, microscopy). A key advantage of our method is that we do not tradeoff any of accuracy, speed, or robustness for the others, thus being a powerful registration algorithm.
Our method shows consistent improvements and robust performance on four community reference brain MRI datasets. Researchers have developed numerous classical image registration methods specifically for neuroimaging studies 13, 17, 41, 42 but registration remains an open challenge in brain mapping 43, 44. FireANTs’ consistent improvement in performance can be attributed to the quasi-second-order update, which normalizes the varying curvature of per-pixel gradients, leading to faster convergence and better local minima. This performance is consistent across metrics (Fig. 1) and anatomical structures (Fig. 10). With the acquisition of larger datasets 45 and high-resolution imaging 27, 46, 32 owing to rapidly growing image acquisition technology and innovations, fast runtime and accurate registration become imperative to enable large-scale studies. Our performance comes with a reduction of runtime of up to three orders of magnitude on standard datasets.
We also demonstrate competitive performance in the EMPIRE10 challenge, widely regarded as a comprehensive evaluation of registration algorithms 47, 48 for lung CT images. Unlike the brain imaging datasets, the EMPIRE10 dataset contains images with large deformations, anisotropic image spacings and voxel sizes, thin structures like airways and pulmonary fissures, which are hard to align based on image intensity alone. These image volumes are typically much larger than what deep learning methods can currently handle at native resolution 48, 39, 30. FireANTs performs much better registration in terms of landmark, fissure alignment, and singularities, while being two orders of magnitude faster than existing state-of-the-art. This experiment also calls attention to a much-overlooked detail - the performance gap due to the choice of representation of diffeomorphisms (direct optimization versus exponential map). We show that direct optimization of the group element is preferable to exponential maps, both in FireANTs and in baselines (ANTs versus DARTEL). This improvement can be attributed to the representation - one can interpret direct optimization as integrating a set of time-dependent velocity fields since the gradients change throughout optimization, allowing more flexibility in the space of diffeomorphisms it can represent, whereas SVF performs the integral of a time-independent velocity field by design. Moreover, computing the exponential map is expensive for diffeomorphisms since the number of iterations can be higher for larger deformations 17. For example, in Fig. 4(a), the exponential map representation (DARTEL) takes substantially longer to run than ANTs. Shooting methods modify the velocity field at each iteration and tend to be sensitive to hyper-parameter choices. For example, in Fig. 2, the results for shooting methods are substantially worse than for methods that optimize the transformation directly. We also observe this for the LPBA40 dataset in Fig. 15, where the shooting method consistently underperformed over a wide range of hyperparameter choices.
FireANTs is consistently 300–2000$\times$ faster than existing state-of-the-art, is robust to the choice of hyperparameters, allowing users to utilize principled hyperparameter search algorithms for novel applications or modalities. This paradigm shift in runtime will enable fast and accurate registration of high-resolution mesoscale and microscale imaging data that will play a paramount role in advancing our understanding of connectomics, neuroscience, cellular and molecular biology, genetics, pathology, among many other disciplines in the biomedical and biological sciences. With breakthrough advances in high-resolution, high-throughput imaging techniques, it is imperative for registration algorithms to also scale with the inordinate amounts of data.
In summary, FireANTs is a powerful and general-purpose multi-scale registration algorithm and sets a new state-of-the-art benchmark. We propose to leverage the accurate, robust, and fast library to speed up registration workflows for ever-growing needs of performant and fast image registration in a spectrum of disciplines within the biomedical and biological sciences, wherein algorithms are bottlenecked by scalability.
4 Methods
4.1 Preliminaries
Given $d$-dimensional images $I:\Omega\to\mathbb{R}^{d}$ and $I^{\prime}:\Omega\to\mathbb{R}^{d}$ where the domain $\Omega$ is a compact subset of $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$, image registration is formulated as an optimization problem to find a transformation $\varphi$ that warps $I^{\prime}$ to $I$. The transformation can belong to an algebraic group, say $G$, whose elements $g\in G$ act on the image by transforming the domain as $(I\circ g)(x)=I(g(x))$ for all $x\in\Omega$. The registration problem solves for
$\varphi^{*}=\operatorname*{\text{argmin}}_{\varphi\in G}L(\varphi)\doteq C(I,I^{\prime}\circ\varphi)+R(\varphi)$ (5)
where $C$ is a cost function, e.g., that matches the pixel intensities of the warped image with those of the fixed image, or local normalized cross-correlation or mutual information of image patches. There are many types of regularizers $R$ used in practice, e.g., total variation, elastic regularization 49, enforcing the transformation to be invertible 50, or volume-preserving 51 using constraints on the determinant of the Jacobian of $\varphi$, etc. If, in addition to the pixel intensities, one also has access to label maps or different anatomical regions marked with correspondences across the two images, the cost $C$ can be modified to ensure that $\varphi$ transforms these label maps or landmarks appropriately.
4.1.1 Properties of the considered Transformation Group
A diffeomorphism is a smooth and invertible map with a corresponding differentiable inverse map 52, 53, 54. We denote the set of all diffeomorphisms on $\Omega$ as $\text{Diff}(\Omega;\mathbb{R}^{d})$. It is useful to note that unlike rigid or affine transforms that have a fixed number of parameters, diffeomorphisms require dense and variable parameterization, typically proportional to the size of the image. When groups of transformations on continuous domains are endowed with a differentiable structure, they are called Lie groups. Diffeomorphisms are also examples of Riemannian manifolds, and are amenable to Riemannian optimization (see  Section 4.3.2).
In this work, we only consider a subgroup of diffeomorphisms. Consider the set of continuously differentiable functions $u\in C^{1}_{0}(\Omega,\mathbb{R}^{d})$ such that $u,J(u)=0$ on $\partial\Omega$, where $J(u)$ is the Jacobian of $u$, such that $[J(u)(x)]_{ij}=\frac{\partial u(x)_{i}}{\partial x_{j}}$. These functions can be extended to have $u\equiv 0$ outside $\Omega$. Then, for a small enough $\epsilon>0$, $x+\epsilon u(x)$ is a diffeomorphism (Proof in Proposition 8.6 in  54). Although these diffeomorphisms are close to identity, diffeomorphisms with larger deviations from the identity can be constructed by composing these ‘small diffeomorphisms’. Therefore, we study the subgroup of diffeomorphisms of the form
$\phi_{n}=(id+\epsilon_{1}u_{1})\circ\ldots\circ(id+\epsilon_{n}u_{n})$ (6)
where $u_{i}$s are defined as before. We denote this subgroup as $G(\Omega,\mathbb{R}^{d})$. This subgroup retains the group structure with identity element $id$, the composition operation $\circ$ induced from $\text{Diff}(\Omega,\mathbb{R}^{d})$, and the inverse group element: $\phi_{n}^{(-1)}=(id+\epsilon_{n}u_{n})^{(-1)}\circ\ldots\circ(id+\epsilon_{1}u_{1})^{(-1)}$ (as each individual $id+\epsilon_{n}u_{n}$ is shown to have an inverse 54). The elements of this subgroup can be thought of as diffeomorphisms arising from time-varying continuously differentiable flows.
However, the rate of convergence of these algorithms are contingent on the severity of ill-conditioning of  Eq. 5. In subsequent sections, we first show the extent of ill-conditioning for diffeomorphic registration which subsequently warrants adaptive optimization over this subgroup of diffeomorphisms.
4.2 Deformable Image Registration is a severely ill-conditioned problem
The ill-conditioned nature of image registration represents a comparatively neglected domain of inquiry within the extant literature. Recent works in the literature  55, 56 only speculate the ill-conditioned nature of registration but do not quantify it. Computing the ill-conditioning requires us to analyze the Hessian of the registration cost function. This is infeasible in general due to the high dimensionality of the problem; the full Hessian of a MRI brain registration problem requires more than 15 petabytes of memory to store. However, we consider a typical scenario of T1-weighted 3D MRI image registration with the L2 loss 30, 15, 16: i.e. $C(I_{f},I_{m},\varphi)=\sum_{i}(I_{f}(\mathbf{x}_{i})-I_{m}(\varphi(\mathbf{x}_{i})))^{2}$. In this case, the gradient of $C$ w.r.t. $\varphi(\mathbf{x}_{i})$ is $(I_{m}(\varphi(\mathbf{x}_{i}))-I_{f}(\mathbf{x}_{i}))\nabla I_{m}(\varphi(\mathbf{x}_{i}))$, which does not depend on $\varphi(\mathbf{x}_{j}),j\neq i$. Therefore, the full Hessian is simply a block-diagonal matrix containing pixelwise Hessians $H_{i}=\nabla^{2}_{\varphi(\mathbf{x}_{i})}C$ with eigenvalues $\{\lambda_{i};i=\{1,2,3\}\}$. This makes the conditioning analysis tractable. We calculate the per-pixel condition number, defined as $\kappa_{i}=|\lambda_{i}^{\text{max}}|/|\lambda_{i}^{\text{min}}|$; and investigate the relationship between the fraction of foreground pixels and $\kappa_{i}$ across multiple spatial resolutions of the images. The study considers three downsampling factors: 1x (original resolution), 2x, and 4x, in accordance with existing multi-scale optimization techniques. Fig. 6 shows that across all resolutions, more than 60% of foreground pixels have a condition number greater than 10. To elucidate the impact of poor conditioning on optimization, we construct a simplified example of an ill-conditioned two-dimensional convex optimization problem, detailed in Section A.3. Even with $\kappa=10$, convergence slows down drastically for an ill-conditioned convex optimization problem. This indicates severe ill-conditioning of the registration problem, strongly motivating the need for first-order adaptive optimization.
Figure 6: Deformable image registration is ill-conditioned. To quantitatively examine ill-conditioning in registration, we compute the distribution of per-pixel condition number for a MRI registration task, at different image downsampling factors (denoted as 1x, 2x, and 4x). A high condition number signifies exacerbated ill conditioning and requires higher-order optimization. A horizontal dashed line denoting $\kappa=10$ is drawn as a reference for substantial ill conditioning. Across all scales, a substantial fraction of foreground voxels are ill-conditioned ($\kappa>10$), necessitating adaptive first-order optimization for faster convergence and accurate registration.
4.3 Adaptive Optimization for Diffeomorphisms
We provide a brief overview of the mathematical frameworks employed to optimize parameters that reside on Riemannian manifolds like diffeomorphisms, followed by a novel algorithm that exploits the group action to define a gradient descent algorithm that eliminates computationally expensive steps. This novel formulation of the ‘gradient descent’ algorithm can then be formulated to incorporate adaptive algorithms such as Adam 26 to optimize diffeomorphisms.
4.3.1 Euclidean gradient descent using the Lie algebra in shooting methods
Each Lie group has a corresponding Lie algebra $\mathfrak{g}$ which is the tangent space at identity. This creates a locally one-to-one correspondence between elements of the group $g\in G$ and elements of its Lie algebra $v\in\mathfrak{g}$ given by the exponential map $\exp:\mathfrak{g}\to G;$ effectively to reach $g=\exp(v)\text{id}$ from identity $\text{id}\in G$, the exponential map dictates that the group element has to move along $v$ for unit time along the manifold. Exponential maps for many groups can be computed analytically, e.g., Rodrigues transformation for rotations, Jordan-Chevalley decomposition 57, or the Cayley Hamilton theorem 58 for matrices. For diffeomorphisms, the Lie algebra is the space of all smooth velocity fields $v:\Omega\to\mathbb{R}^{d}$. There exist iterative methods to approximate the exponential map called the scaling-and-squaring approach 17, 30 which uses the identity
$\varphi=\exp(v)=\lim_{N\to\infty}\left(\text{id}+\frac{v}{N}\right)^{N}$
to define a recursion by choosing $N$ to be a large power of 2, i.e. $N=2^{M}$ as
$\varphi^{(1/2^{M})}$ $=id+v/2^{M}$
$\varphi^{(1/2^{k})}$ $=\varphi^{(1/2^{(k+1)})}\circ\varphi^{(1/2^{(k+1)})}\quad\quad\forall k\in\{0,1\ldots,M-1\};$
This can be thought of as a special case of  Eq. 6 with $n=2^{M},\epsilon=\frac{1}{n}$ and $u_{1}=\ldots=u_{n}=v$.
By virtue of the exponential map, we can solve the registration problem of finding $\varphi\in G$ by directly optimizing over the Lie algebra $v$. This is because the Lie algebra is a vector space and we can perform, for example, standard Euclidean gradient descent for registration 59, 60, 61. Such methods are called stationary velocity field or shooting methods. At each iteration, one uses the exponential map to get the transformation $\varphi$ from the velocity field $v$, computes the gradient of the registration objective with respect to $\varphi$, pulls back this gradient into the tangent space where $v$ lies
$\nabla_{v}L=\frac{\partial\varphi}{\partial v}\nabla_{\varphi}L$
and finally makes an update to $v$. Traditional methods like DARTEL 17 implement this approach. This is also very commonly used by deep learning methods for registration 30, 62, 63 due to its simplicity. Geodesic shooting methods are more sophisticated implementations of this approach where $\varphi$ is the solution of a time-dependent velocity that follows the geodesic equation; the geodesic is completely determined by the initial velocity $v_{0}\in\mathfrak{g}$.
Adaptive optimization algorithms can be applied to the Lie algebra since it a Euclidean vector space $\mathfrak{g}$. However, there a number of challenges with this method. First, this method requires computing the exponential map and its derivative, both of which need to be iteratively evaluated at each step of gradient descent. This is evident in  Fig. 2 where direct optimization with ANTs runs much faster than the Lie-algebra counterpart. Moreover, the exponential map is only locally diffeomorphic, meaning it is suitable for modelling deviations close to the identity but not for large deformations – this leads to less expressivity and poor performance. In Fig. 2, the greedy SyN method which employs direct optimization significantly outperforms the Lie algebra-based DARTEL. In Fig. 15 we observed that across a large variety of hyper-parameters evaluated via grid search, direct optimization consistently led to better target overlap compared to its Lie algebra counterpart on the LPBA40 dataset. Therefore, we do not consider this method in our work.
4.3.2 Riemannian gradient descent
Solving the registration problem directly on the space of diffeomorphisms avoids repeated computations to and fro via the exponential map. The downside however is that one now has to explicitly account for the curvature and tangent spaces of the manifold. The updates for Riemannian gradient descent 64 at the $t^{\text{th}}$ iteration are
$\varphi_{t+1}$ $=\exp_{\varphi_{t}}\left(-\eta\ \text{Proj}_{\varphi_{t}}(\nabla_{\varphi}L)\right)$ (7)
where $\text{where }\nabla_{\varphi}L$ $=\mathbf{g}^{-1}_{\varphi_{t}}\frac{\partial L}{\partial\varphi},$
where one pulls back the Euclidean gradient $\frac{\partial L}{\partial\varphi}$ onto the manifold using the inverse metric tensor $\mathbf{g}$ (which makes the gradient invariant to the parameterization of the manifold of diffeomorphisms) before projecting it to the tangent space using $\text{Proj}_{\varphi_{t}}$. Since the tangent space is a local first-order approximation of the manifold’s surface, we can move along this descent direction by a step-size $\eta$ and compute the updated diffeomorphism $\varphi_{t+1}$, represented as the exponential map from $\varphi_{t}$ computed in the direction of $-\text{Proj}_{\varphi_{t}}(\nabla_{\varphi}L)$.
However, there are a few challenges in optimizing diffeomorphisms using Riemannian gradient descent. First, adaptive optimization algorithms such as RMSProp 25, Adagrad 65 and Adam 26 have become popular because they can handle poorly conditioned optimization problems in deep learning. Variants for optimization on low-dimensional Riemannian manifold exist 66, 67, 68, 69. In contrast to these manifolds, diffeomorphisms are a high-dimensional variable-sized group (e.g., the parameterization of the warp field scales with that of the image size). Therefore, operations like computing the Riemannian metric tensor, and parallel transport of the optimization state variables (momentum and curvature) are very computationally expensive. For diffeomorphisms, computing the parallel transport requires solving a system of partial differential equations, which is computationally expensive. For these reasons, we do not consider direct Riemannian optimization for diffeomorphisms in our work.
4.4 Exploiting the group structure of diffeomorphisms
Diffeomorphisms are imbued with additional structure compared to a Riemannian manifold – they are a Lie group as well. Not all Riemannian manifolds are Lie groups - notable examples of non-Lie group Riemannanian manifolds include the sphere $\mathbb{S}^{n}$, fixed-rank matrices, and the Stiefel and Oblique manifolds 64. The additional Lie group structure of $G(\Omega,\mathbb{R}^{d})$ allows us to exploit the group action to define a gradient descent algorithm that eliminates computationally expensive steps. In the following text, we provide a novel method to compute a descent direction in the group of diffeomorphisms that is computationally efficient and can be used with adaptive optimization algorithms.
Minimizing the Eulerian differential.
Consider a function $U:G\rightarrow\mathbb{R}$ that we aim to minimize. Let $V$ be an admissible Hilbert space of vector fields on $\Omega$ embedded in $C^{1}_{0}(\Omega,\mathbb{R}^{d})$. We define an Eulerian differential in $V$ if there exists a linear form $\partial\bar{U}\in V^{*}$ such that for all $v\in V$:
$\left(\partial\bar{U}(\varphi)|v\right)_{E}=\partial_{t}U(\varphi\circ\varphi^{v}_{0t})\Big{|}_{t=0}$ (8)
This definition of Eulerian differential is different from the one in  54 to perform all updates ($v$) in the tangent space at identity and leverage Jacobian-free descent (see later). The goal is to choose a suitable $v$ such that the directional change of the Eulerian differential along $v$ is negative, making $v$ a descent direction. A more familiar rate of change of $U$ along a curve $v$ is given by the Gateaux derivative:
$\left(\frac{\partial U}{\partial\varphi}\Bigg{|}v\right)_{G}=\partial_{t}U(\varphi+tv)\Big{|}_{t=0}$ (9)
The Eulerian differential is closely related to the Gateaux derivative of $U$ at $\varphi$ as:
$\left(\partial\bar{U}(\varphi)|v\right)_{E}=\left(\frac{\partial U}{\partial\varphi}\Bigg{|}J(\varphi)v\right)_{G}$
using chain rule. The right side is further expanded as:
[…]
where $J(\varphi)(x)=J(\varphi(x))$ with slight abuse of notation. We introduce the Gateaux derivative and relate it to the Eulerian derivative because we typically have access to the Gateaux derivative using automatic differentiation tools like PyTorch, but to perform optimization on the group of diffeomorphisms, we need to compute the Eulerian differential.
Choosing
$v_{d}(x)=-J(\varphi(x))^{T}\frac{\partial U}{\partial\varphi}(\varphi)(x)$
gives us:
[…]
This choice of $v_{d}(x)$ is therefore a descent direction for the Eulerian differential of $U$ at $\varphi$. To perform gradient descent on the Eulerian differential at $\varphi$, we need to compute the descent direction $v_{d}$, perform the exponential map with a small learning rate $\eta_{t}$, and perform the update:
$\varphi_{t+1}=\varphi_{t}\circ\exp_{id}(\eta_{t}v_{d})$
For small enough $\eta_{t}$, the exponential map can be approximated with a retraction map (i.e. $\exp_{id}(\eta_{t}v_{d})\approx id+\eta_{t}v_{d}$), which is quick to compute.
We quickly contextualize the key differences between Gateaux gradient descent and our proposed Eulerian descent. First, the steepest descent direction in Gateaux gradient descent is $-\frac{\partial U}{\partial\varphi}(\varphi)$, whereas it is $-J(\varphi)^{T}\frac{\partial U}{\partial\varphi}(\varphi)$ in Eulerian descent. Second, the update rule in Gateaux gradient descent is $\varphi_{t+1}=\varphi_{t}-\eta_{t}\frac{\partial U}{\partial\varphi}(\varphi)$, whereas it is $\varphi_{t+1}=\varphi_{t}\circ\exp_{id}(\eta_{t}v_{d})$ in Eulerian descent. These two differences capture the essence of performing optimization on the group of diffeomorphisms in contrast to optimizing on the (Euclidean) ambient space directly.
Adaptive optimization on diffeomorphisms.
Note that for small enough $t$, the descent direction $v_{d}(x)$ can also be interpreted as a vector in the tangent space at identity, with $\varphi_{0t}^{v}=\exp_{id}(tv)$ since $\varphi_{00}^{v}=id$, and $\partial_{t}\varphi_{0t}^{v}|_{t=0}=v$. Descent directions over gradient descent iterations $i$ denoted as $v_{d}^{(i)}$ all lie on the same vector space, i.e. the tangent space at identity. Therefore, first order algorithms like Adam can be applied on the sequence of descent directions $v_{d}^{(i)}$ which now lie in the same vector space, without requiring computing the metric tensor, parallel transport or change of coordinates (charts) throughout the optimization process. This framework leveraging the group structure forms the core of our adaptive optimization algorithm for diffeomorphisms. Our framework is therefore a significant advantage over Riemannian optimization methods which require parallel transport of the momentum and curvature vectors at each iteration.
4.4.1 Jacobian-Free Eulerian Descent
We provided an obvious choice of descent direction $v_{d}(x)$ for the Eulerian differential of $U$ at $\varphi$. The Gateaux derivative $\frac{\partial U}{\partial\varphi}$ is readily obtained using automatic differentiation tools like PyTorch. However, the descent direction requires us to multiply this derivative with the Jacobian of the diffeomorphism $J(\varphi)$, which may be computationally expensive. However, in most diffeomorphic image registration applications, the role of the diffeomorphism is warp the image by performing local translations, scaling and shearing without introducing large local rotations. Mathematically, we consider the polar form of the Jacobian $J(\varphi)(x)=U(x)P(x)$ where $U(x)$ is a unitary matrix, and $P(x)$ is a positive definite matrix. We assume that for most applications, $U(x)\approx I_{d\times d}$, making $J(\varphi)(x)$ positive definite. With this assumption, we can choose the modified descent direction
$v_{d}^{\prime}(x)=-\frac{\partial U}{\partial\varphi}(\varphi)(x)$
and the Eulerian differential at $\varphi$ is
[…]
since $v_{d}^{\prime}(x)^{T}J(\varphi(x))v_{d}^{\prime}(x)\geq 0$ for all $x\in\Omega$, owing to the (assumed) positive definiteness of $J(\varphi)(x)$. For all experiments, Jacobian-free descent directions $v_{d}^{\prime}(x)$ are used, and they provide faster runtime and with same accuracy. Adaptive first-order optimization can now be performed on this modified sequence on descent directions $v_{d}^{\prime(i)}(x)$, saving significant computational and memory overhead by avoiding computation of $J(\varphi)$.
Note that this algorithm using the Eulerian differential is only possible due to the group structure of diffeomorphisms. For an arbitrary Riemannanian manifold $\mathcal{M}$ and points $\varphi,\varphi_{0t}^{v}\in\mathcal{M}$, the operation $\varphi\circ\varphi_{0t}^{v}$ does not make sense. The additional group structure of $G(\Omega,\mathbb{R}^{d})$ allows us to propose a novel Eulerian descent algorithm without performing Lie algebra optimization, or Riemannian gradient descent, both of which are computationally expensive for diffeomorphisms.
4.5 Interpolation strategies for multi-scale registration
Classical approaches to deformable image registration is performed in a multi-scale manner. Specifically, an image pyramid is constructed from the fixed and moving images by downsampling them at different scales, usually in increasing powers of two. Optimization is performed at the coarsest scale first, and the resulting transformation at each level is used to initialize the optimization at the next finer scale. Specifically, for the fixed image $I$ and the moving image $I^{\prime}$ and $K$ levels, let the downsampled versions be $\{I_{k}\}_{k=1}^{K}$ and $\{I^{\prime}_{k}\}_{k=1}^{K}$, where $k$ is the scale index from coarsest to finest. At the $k$-th scale, the transformation $\varphi_{k}$ is optimized as
$\varphi_{k}^{*}=\operatorname*{\text{argmin}}_{\varphi_{k}\in G}L(I_{k},I^{\prime}_{k}\circ\varphi_{k})$
where $\varphi_{k}$ is initialized as
$\varphi_{k}=\begin{cases}{id}&\text{if }k=1\\ \text{Upsample}(\varphi_{k-1})&\text{otherwise}\end{cases}$
Unlike existing gradient descent based approaches, our Riemannian adaptive optimizer also contains state variables $m_{k}$ corresponding to the momentum and $\nu_{k}$ corresponding to the EMA of squared gradient, at the same scale as $\varphi_{k}$, which require upsampling as well.
Unlike upsampling images, upsampling warp fields and their corresponding optimizer state variables requires careful consideration of the interpolation strategy. Bicubic interpolation is a commonly used strategy for upsampling images to preserve smoothness and avoid aliasing. However, bicubic interpolation of the warp field can lead to overshooting, leading to introducing singularities in the upsampled displacement field when there existed none in the original displacement field. In contrast, bilinear or trilinear interpolation does not lead to overshooting, and therefore diffeomorphism of the upsampled displacement is guaranteed, if the original displacement is diffeomorphic. We demonstrate this using a simple 2D warp field in  Fig. 12(b). On the left, we consider a warp field created by nonlinear shear forces. This warp field does not contain any tears or folds - and is diffeomorphic. We upsample this warp field using bicubic interpolation (top) and bilinear interpolation (bottom). We also plot a heatmap of the negative of the determinant of the Jacobian of the upsampled warp, with a white contour representing the zero level set. Qualitatively, bicubic interpolation introduces noticable folds in the warping field, leading to non-diffeomorphisms in the upsampled warp field. The heatmap shows a significant portion of the upsampled warp field has a negative determinant, indicating non-invertibility. On the other hand, bilinear interpolation looks blocky but preserves diffeomorphism everywhere, as also quantitatively verified by the absence of a zero level set in the heatmap.
The complete algorithm is described in  Algorithm 1.
Algorithm 1 FireANTs
1:Input: Fixed image $I_{f}$, Moving image $I_{m}$
2:Scales $[s_{1},s_{2},\ldots,s_{n}]$, Iterations $[T_{1},T_{2},\ldots T_{n}]$, $n$ scales
3:optstate optimizer state (for Adam, RMSProp, etc.)
4:use_jac boolean specifying whether to use Jacobian in descent direction
5:
6:Initialize $\varphi\leftarrow\mathbf{id}_{s_{1}}$. $\triangleright$ Initialize warp to identity at first scale
7:Initialize ${l}\leftarrow 1$. $\triangleright$ Initialize current scale
8:while ${l}\leq n$ do
9:     Initialize $i\leftarrow 0$
10:     Initialize $I_{f}^{l},I_{m}^{l}\leftarrow\text{downsample}(I_{f},s_{l}),\text{downsample}(I_{m},s_{l})$
11:     while $i<T_{{l}}$ do
12:         $U_{i}\leftarrow C(I_{f}^{l},I_{m}^{l}\circ\varphi^{i})+R(\varphi)$
13:         Compute $v_{d}^{\prime}(x)\leftarrow\frac{\partial U_{i}}{\partial\varphi}(\varphi^{(i)})(x)$triangleright Jacobian-free Eulerian descent direction
14:         if use_jac then
15:              Compute $v_{d}^{\prime}(x)\leftarrow J^{T}(\varphi^{(i)}(x))v_{d}^{\prime}(x)$triangleright Eulerian descent direction
16:         end if
17:         Update $(v_{d}^{\prime}(x),\text{{optstate}})\leftarrow\text{{optstate}}(v_{d}^{\prime}(x))$triangleright Apply and update optimizer state
18:         Update $\varphi^{(i+1)}\leftarrow\varphi^{(i)}\circ\exp_{id}(\epsilon_{i}v_{d}^{\prime})\approx\varphi^{(i)}\circ(id+\epsilon_{i}v_{d}^{\prime})$
19:         $i\leftarrow i+1$
20:     end while
21:     if $l<n$ then
22:         $\varphi\leftarrow\text{Upsample}(\varphi,s_{(l+1)})$triangleright Upsample warp to next scale using bilinear/trilinear interpolation
23:     end if
24:     $l\leftarrow l+1$
25:end while
Figure 7: Algorithm for FireANTs Algorithm 1 outlines the key steps in FireANTs - computing the Jacobian-free Eulerian descent direction which is simply the Gateaux derivative. If the boolean use_jac is specified, then use the steepest Eulerian descent direction instead. This descent direction is then modified using any adaptive optimization algorithm denoted as optstate. The warp field is then updated using the exponential map or retraction map for small $\epsilon_{i}$. After optimization at a given scale, the warp field is upsampled using bilinear or trilinear interpolation to the next scale until optimization is complete for all steps.
Appendix A Supplementary Information
A.1 Datasets used in Klein et al.challenge
Brain image data and their corresponding labels for 80 normal subjects were acquired from four different datasets. The LPBA40 dataset contains 40 brain images and their labels to construct the LONI Probabilistic Brain Atlas (LPBA40). All volumes were skull-stripped, and aligned to the MNI305 atlas  70 using rigid-body transformation to correct for head tilt. For all these subjects, 56 structures were manually labelled and bias-corrected using the BrainSuite software. The IBSR18 dataset contains brain images acquired at different laboraties through the Internet Brain Segmentation Repository. The T1-weighted images were rotated to be in Talairach alignment and bias-corrected. Manual labelling is performed resulting in 84 labeled regions. For the CUMC12 dataset, 12 subjects were scanned at Columbia University Medical Center on a 1.5T GE scanner. Images were resliced, rotated, segmented and manually labeled, leading to 128 labeled regions. Finally, the MGH10 dataset contains 10 subjects who were scanned at the MGH/MIT/HMS Athinoula A. Martinos Center using a 3T Siemens scanner. The data is bias-corrected, affine-registered to the MNI152 template, and segmented. Finally the images were manually labeled, leading to 74 labeled regions. All datasets have a volume of $256\times 256\times\{128,124\}$ voxels with varying amounts of anisotropic voxel spacing, ranging from $0.84\times 0.84\times 1.5$mm to $1\times 1\times 1.33$mm.
A.2 Modular software implementation to enable effective experimentation
Registration is a key part of many data processing pipelines in the clinical literature. Our software implementation is designed to be extremely flexible, e.g., it implements a number of existing registration methods using our techniques, modular, e.g., the user can choose different group representations (rigid or affine transforms, diffeomorphisms), objective functions, optimization algorithms, loss functions, and regularizers. Users can also stack the same class of transformations, but with different cost functions. For example, they can fit an affine transform using label maps and Dice loss, and use the resultant affine matrix as initialization to fit another affine transform using the cross-correlation registration objective. This enables seamless tinkering and real-time investigation of the data. Deformations can also be composed in increasing order of complexity (rigid $\to$ affine $\to$ diffeomorphisms), thereby avoiding multiple resampling and subsequent resampling artifacts. We have developed a simple interface to implement custom cost functions, which may be required for different problem domains, with ease; these custom cost functions can be used for any of the registration algorithms out-of-the-box. Our implementation can handle images of different sizes, anisotropic spacing, without the need for resampling into a consistent physical spacing or voxel sizes. All algorithms also support multi-scale optimization (even with fractional scales) and convergence monitors for early-stopping.
Our software is implemented completely using default primitives in PyTorch. All code and example scripts is available at
A.3 On the Ill-conditioning of Image Registration
Image registration is a highly ill-conditioned, and non-convex problem necessitating advanced optimization methods for convergence. To provide more intuition on the effect of $\kappa$ on convergence of the SGD algorithm, we consider a toy example of a 2D optimization problem. Specfically, we consider a loss function $f_{\kappa}(x,y)=x^{2}+\kappa y^{2}$ where $\kappa>1$ becomes the condition number of the problem. Qualitatively, the effect of the first term diminishes exponentially fast with $\kappa$ (Fig. 8(a)). Quantitatively, we run both SGD and Adam optimization for a 1000 iterations starting from the point $(x,y)=(5,5)$. Fig. 8(c) shows that SGD works extremely well for $\kappa=1$ which is the best-conditioned loss function, but quickly gets stuck for $\kappa\geq 100$. On the contrary, Adam is invariant to the condition number and converges to the minima for all values of $\kappa$. This is because for a diagonal Hessian (as in this case), the second-order adaptive terms are proportional to the diagonal elements of the Hessian. These condition numbers are vanishingly small compared to those in typical image registration tasks, which can exceed $10^{5}$, making them extremely ill-conditioned.
We also consider a more realistic, but tractable scenario of the convex loss function $f_{\kappa,\theta}(x,y)=x_{\theta}^{2}+\kappa y_{\theta}^{2}$, where
$\begin{bmatrix}x_{\theta}\\ y_{\theta}\end{bmatrix}=\begin{bmatrix}\cos(\theta)&\sin(\theta)\\ -\sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}$
We choose $\theta=\pi/3$ for this experiment. This is simply a rotated version of the previous family of loss functions, as shown in Fig. 8(b). The trajectories obtained from optimization using SGD (Fig. 8(d)) are virtually identical to that in Fig. 8(c) since the new gradients are simply rotated versions of the previous gradients, and the distance from the minima is invariant to the rotation. However, the trajectories from Adam optimization are qualitatively very different, owing to the increasing difference between the true Hessian and its diagonal approximation. Even so, the final point is at a distance of less than $10^{-3}$ units to the minima for $\kappa=1000$, showing the effectiveness of adaptive optimization even for ill-conditioned, non-diagonal Hessians. This is a strong motivation to extend adaptive optimization for non-Euclidean diffeomorphic registration, which is very high-dimensional and ill-conditioned.
(a) Log-Loss landscape of the toy problem $f_{\kappa}(x,y)=x^{2}+\kappa y^{2}$ for $\kappa=1,10,100,1000$. The log-loss becomes increasingly sharp along the y-direction as $\kappa$ increases.
(b) Log-Loss landscape of the toy problem $f_{\kappa}(x,y)=x_{\theta}^{2}+\kappa y_{\theta}^{2}$ for $\kappa=1,10,100,1000$, where $(x_{\theta},y_{\theta})$ is the coordinate $(x,y)$ rotated by an angle $\theta$ about the origin.
(c) Optimization of Fig. 8(a) using SGD and Adam shows that SGD fails to recover the minima for $\kappa\geq 100$ while Adam is invariant to the condition number for diagonal Hessian matrices. This is a strong motivation to use first order adaptive optimization for registration where the condition number can exceed $10^{5}$.
(d) Optimization of Fig. 8(b) using SGD shows identical optimization trajectories as Fig. 8(c). Adam, however, is not invariant to the condition number because the difference between the true Hessian and its diagonal approximation increases with $\kappa$. Even so, the final point is at a distance of less than $10^{-3}$ units to the minima, showing the mitigating effect of adaptive optimization even for non-diagonal Hessians.
Figure 9: Overview of FireANTs and its features: (a) shows the modalities our method is tested on. We demonstrate results on in-vivo T1-weighted brain MRI, lung CT, and expansion microscopy volumes. FireANTs can optimize intensity images as well as binary masks (lung masks in CT) or entire anatomical label maps (brain MRI). (b) shows the technical contributions of FireANTs. We extend Adaptive Optimization to multi-scale Diffeomorphisms by first writing the Riemannian gradient update, and then avoiding parallel transport of the optimization state by leverging the interchangability of the Riemannian gradient at arbitrary transform $\varphi_{t}$ with the Riemannian gradient at $\varphi=\textbf{Id}$. For the Lie-algebra representation, the Gateaux derivative $\frac{\partial L}{\partial\varphi}$ is projected to $\frac{\partial L}{\partial v}$ using analytical backprop. Since the Lie algebra is a vector space, we use standard adaptive optimizers (see Section 4.3.2 for more details). (c) takes a closer look at multi-scale interpolation for diffeomorphisms represented as a warp field. Bicubic interpolation can introduce folding of the warp field at a finer resolution due to overshooting, but bilinear interpolation does not. Therefore, we use this for interpolating the warp field and the optimizer state. (d) shows the extensive experimental setup. Our method is orders of magnitude faster, has state-of-the-art performance on 3 challenges, is robust across hyperparameters and datasets, and is modular and easy to extend.
Figure 10: Regionwise target overlap on the brain MRI datasets: We further evaluate regionwise overlap scores by sampling 15 regions from each dataset, and comparing their distribution using our method and ANTs. Our method has a much higher median score, and better interquartile ranges across regions, demonstrating both accuracy and robustness.
Figure 11: Comparison of our method with ANTs on 4 MRI brain datasets: Registration quality is validated by measuring volume overlap of label maps between the fixed and warped label maps. (a): For anatomical region $r$, warped (binary) label map ${S}_{r}$ and fixed label map ${T}_{r}$, target and mean overlap are defined as $|{S}_{r}\cap{T}_{r}|/|{T}_{r}|$ and $2|{S}_{r}\cap{T}_{r}|/(|{S}_{r}|+|{T}_{r}|)$. We define the aggregate target overlap over all anatomical regions as $\sum_{r}(|{S}_{r}\cap{T}_{r}|/|{T}_{r}|)$ and Klein et al. 29 define it as $(\sum_{r}|{S}_{r}\cap{T}_{r}|)/(\sum_{r}|{T}_{r}|)$, likewise for other metrics. The latter aggregation is denoted with the suffix (Klein) in the figure. In all four datasets, the boxplots show a narrower interquartile range and substantially higher median than ANTs (higher is better), underscoring the stability and accuracy of our algorithm. (b): Other measures of anatomical label overlap used in 29 are false positives ($|{T}_{r}\backslash{S}_{r}|/|{T}_{r}|$), false negatives ($|{S}_{r}\backslash{T}_{r}|/|{S}_{r}|$), and volume similarity ($2(|{S}_{r}|-|{T}_{r}|)/(|{S}_{r}|+|{T}_{r}|)$) (lower is better). We observe similar trends as in (a), with a narrower interquartile range and substantially lower median values. Results of per region overlap metrics are in the  Fig. 10.
(a) Trick to avoid parallel transport in Riemannian Adaptive Optimization using Eulerian differentials
(b) Bicubic interpolation of diffeomorphic map does not preserve diffeomorphism
Refer to captionRefer to caption
Figure 12: Overview of tricks for multi-scale adaptive optimization for diffeomorphisms: (a) We exploit the group structure of diffeomorphisms to define an Eulerian differential that avoids the need for parallel transport in adaptive optimization algorithms. (b) We show the effect of downsampling on the warp and determinant of the Jacobian for a single image pair. The first column shows the initial warp, and the second and third columns show the warp and determinant of the Jacobian for the cubic and bilinear interpolations, respectively.
Figure 13: Qualitative results on EMPIRE10 challenge: (a) shows the fixed image, (b) shows the registration performed by ANTs, and (c) our method, all with zoomed in regions. ANTs performs a coarse registration with ease, but still leaves out critical alignment of lung boundary and airways by not utilizing adaptive optimization. Our method performs perfectly diffeomorphic registration by construction, and does not lead to any registration errors, both in the lung boundaries or internal features.
Figure 14: More Qualitative results on EMPIRE10 challenge: (a) shows the fixed image, (b) shows the registration performed by ANTs, and (c) our method, all with zoomed in regions. ANTs performs a coarse registration with ease, but still leaves out critical alignment of lung boundary and airways by not utilizing adaptive optimization. Our method performs perfectly diffeomorphic registration by construction, and does not lead to any registration errors, both in the lung boundaries or internal features.
Figure 15: Comparison of exponential versus direct optimization on LPBA40 dataset: We run the hyperparameter grid search on the LPBA40 dataset using direct Riemannian gradient updates with Adam optimizer (denoted as rgd), and optimizing the velocity field by computing the exponential map to represent the diffeomorphism (denoted as exp) across all the configurations shown in Fig. 5(a). The average target overlap for each configuration is then stored, and a histogram of target overlap values of the dataset is constructed. Note that the rgd variant has a significantly more number of configurations near the optimal value, and the average performance and the overall distribution of our optimization is better for the rgd variant than exp. Similar trends can be observed for the EMPIRE10 lung challenge in Fig. 2, where the exp representation underperforms for the same cost function, data, etc. Therefore, we recommend direct RGD optimization for diffeomorphisms.